# Gradient Flow Finite Element Discretisations with Energy-Based $hp$–Adaptivity for the Gross-Pitaevskii Equation with Angular Momentum Rotation

### v1

@@ 4.1.2. Non-linear GPE with optical lattice potential
In this second example, we consider a more challenging problem with an oscillating potential function $V$; to this end, we select $\Omega=(-6,6)^{2}$, $\beta=1000$, $V=\nicefrac{{1}}{{2}}(x^{2}+y^{2})+20+20\sin(2\pi x)\sin(2\pi y)$, and again $\omega=0$. The computed approximation to the ground-state is depicted in Figure 3(a); in this setting $u_{\mathrm{GS}}$ resembles a ‘hedgehog’ structure consisting of a large number of very thin spikes. Computations undertaken in [28] and [25] give values of the ground-state energy $\mathsf{E}(u_{\mathrm{GS}})\approx 30.40965$ and $\mathsf{E}(u_{\mathrm{GS}})\approx 30.387533$, respectively. Based on exploiting $hp$–adaptivity, we have computed a reference value of $\mathsf{E}(u_{\mathrm{GS}})\approx\underline{30.3874145736}32611$.
The performance of the proposed $hp$–refinement algorithm is depicted in Figure 3(b); here, again we observe exponential convergence of the error in the computed ground-state energy, despite the fine scale structures present in the ground-state. Moreover, $hp$–refinement clearly outperforms its $h$–version counterpart. The final $hp$–mesh shown in Figure 3(c) indicates that some $h$–refinement is necessary across most of the domain, with again $p$–enrichment being undertaken in smooth regions.
@@
Figure 3. Non-linear GPE with optical lattice potential. (a) Computed ground-state $u_{N}$; (b) Comparison of the error with respect to the square root of the number of degrees of freedom; (c) Final $hp$-mesh.
@@ 4.1.3. Non-linear energy functional with a nonsymmetric potential $V$
Figure 4(b) again highlights the potential benefits of employing $hp$–adaptivity: exponential convergence of the error in the computed ground-state energy is observed, as well as its superiority in terms of the computed error versus the number of degrees of freedom employed, when compared with mesh subdivision ($h$–refinement). The final $hp$–mesh depicted in Figure 4(c) highlights that some $h$-refinement is necessary around the top of the hill, while largely $p$–enrichment is undertaken elsewhere.
@@
Figure 4. Non-linear GPE with a nonsymmetric potential $V$. (a) Computed ground-state $u_{N}$; (b) Comparison of the error with respect to the square root of the number of degrees of freedom; (c) Final $hp$-mesh.
@@ 4.2. GPE with angular momentum operator $\omega>0$
Finally, in this section, we now turn our attention to the case when the angular momentum operator is present in the underlying energy functional, i.e., when $\omega>0$. To this end, we set $\Omega=(-L,L)$, for some $L>0$, $V=\nicefrac{{1}}{{2}}(x^{2}+y^{2})$, and consider the range of $\omega$ and $\beta$, for a given $L$, presented in Table 1. In addition, here we report the minimal energy computed using the proposed $hp$–refinement algorithm based on selecting the initial condition
where
Numerical experiments for each of the values of $\omega$ and $\beta$ given in Table 1 are presented in Figures 58. As before, we show the computed ground-state, the comparison between the error in the approximated ground-state energy when employing both $hp$– and $h$–refinement, and the final $hp$–mesh. We note that as $\omega$ is increased, the number of vortices present in the ground-state increases; for the case when $\omega=0.9$ and $\beta=1000$, the ground-state features over 50 vortices arranged in a regular triangular lattice referred to as the Abrikosov lattice. In all cases we again observe that the proposed $hp$–refinement algorithm leads to exponential convergence of the error in the approximated ground-state energy; moreover, the efficiency of employing $hp$–refinement when compared to mesh subdivision alone is clearly highlighted. As the number of vortices increases with each of the test cases considered here, we observe that the accuracy attained for the computed ground-state energy decreases. Moreover, we also observe that more refinement of the computational mesh is necessary, before subsequent $p$–refinement is employed to yield exponential convergence of the error in the ground-state energy, as $\omega$ is increased. In the simplest setting, i.e., when $\omega=0.5$ and $\beta=10$ we observe that the proposed $hp$–refinement algorithm simply defaults to polynomial enrichment only.
@@
Figure 5. GPE with angular momentum: $\omega=0.5$, $\beta=10$. (a) Computed ground-state density $|u_{N}|^{2}$; (b) Comparison of the error with respect to the square root of the number of degrees of freedom; (c) Final $hp$-mesh.
@@
Figure 6. GPE with angular momentum: $\omega=0.75$, $\beta=100$. (a) Computed ground-state density $|u_{N}|^{2}$; (b) Comparison of the error with respect to the square root of the number of degrees of freedom; (c) Final $hp$-mesh.
@@
Figure 7. GPE with angular momentum: $\omega=0.8$, $\beta=500$. (a) Computed ground-state density $|u_{N}|^{2}$; (b) Comparison of the error with respect to the square root of the number of degrees of freedom; (c) Final $hp$-mesh.
@@
Figure 8. GPE with angular momentum: $\omega=0.9$, $\beta=1000$. (a) Computed ground-state density $|u_{N}|^{2}$; (b) Comparison of the error with respect to the square root of the number of degrees of freedom; (c) Final $hp$-mesh.
@@ 5. Conclusions
In this paper, we have considered the stationary non-linear Gross–Pitaevskii eigenvalue problem in the presence of a rotating magnetic field, which yields an intricate topology manifested by the presence of quantized vortices in the ground-state; the number and location of these vortices is unknown a priori. In particular, we have extended the energy-based adaptive finite element strategy presented in  [25] in two key directions: firstly, we have considered the case when a rotating magnetic field is present in the underlying energy functional, and secondly, we have generalised the adaptivity framework to allow for $hp$-refinements of the underlying discrete finite element approximation space. Towards these goals, we have also presented a new result on the convergence of a projected gradient flow (Theorem 2.6) which respects the energy-based topology of the underlying problem at hand. The key feature of the proposed energy-adaptive $hp$-mesh refinement strategy is that it is based solely on potential energy-decay which can be computed locally and does not require sophisticated localisation schemes as is standard for residual-based a posteriori error estimators. The resulting method thereby intertwines the $hp$-mesh refinement iterative procedure with the iterative non-linear solver for the underlying problem within a unified framework in the sense that the numerical approximation to the non-linear problem is computed in tandem while the $hp$-mesh is adaptively refined. The performance of this proposed strategy is demonstrated on a diverse set of benchmark test problems that highlight the exponential convergence of the error in the computed ground-state energy with respect to the number of degrees of freedom in the resulting $hp$-finite element space.
@@ Acknowledgments
T. P. W. acknowledges the financial support of the Swiss National Science Foundation (SNF), Grant No. $212868200021\underline{\phantom{-}}$ 212868. B. S. acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), Project No. 442047500, through the Collaborative Research Center “Sparsity and Singular Structures” (SFB 1481). Finally, we thank Eric Cancès for encouraging discussions and motivating us to extend the framework to the case with rotating magnetic field.
Appendix A Existence of a ground-state
In the following, we verify that there exists at least one ground-state of the energy functional $\mathsf{E}$ from (1) under the assumption (V1). To this end, we first show that $\mathsf{E}:\mathbb{H}\to\mathbb{R}$ is bounded from below and coercive. Indeed, following along the lines of the proof of Proposition 2.2, we obtain that $\mathsf{E}(u)\geq\nicefrac{{c}}{{2}}\left\|\nabla u\right\|_{\mathrm{L}^{2}(\Omega)}^{2},$ with $c>0$ being the constant from (19). Consequently, by the same reasoning as in the proof of [16, Lem. 2], there exists a bounded sequence $\{u_{n}\}_{n}\subset\mathcal{S}_{\mathbb{H}}$ with
$\lim_{n\to\infty}\mathsf{E}(u_{n})=\inf_{v\in\mathcal{S}_{\mathbb{H}}}\mathsf{E}(v),$
from which we can extract a subsequence $\{u_{n_{k}}\}_{k}$ that converges weakly in $\mathbb{H}$ and strongly in $\mathrm{L}^{2}(\Omega)$ to some element $u\in\mathbb{H}$. Due to the convergence in $\mathrm{L}^{2}(\Omega)$, we indeed have that $\left\|u\right\|_{\mathrm{L}^{2}(\Omega)}=1$; i.e., $u\in\mathcal{S}_{\mathbb{H}}$. We further remark that $\mathsf{E}$ is weakly sequentially lower semicontinuous, since it is convex and strongly continuous; see, e.g., [39, Prop. 25.20]. Consequently, we have that
thereby implying that $u$ is indeed a minimiser of the energy functional $\mathsf{E}$.
Appendix B Sketch of proof of Theorem 2.8
We proceed in several steps.
  1. Step 1:
    Firstly, a straightforward calculation reveals that the mass, without the scaling (38a), growths in each iteration step (38b). In particular, we have that
    $\left\|\widehat{u}^{n+1}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}=\left\|u^{n}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|\widehat{u}^{n+1}-u^{n}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}.$ (54)
  2. Step 2:
    It can further be verified that, for $\tau_{n}\leq\nicefrac{{2}}{{5}}$,
    $\begin{split}\mathsf{E}(u^{n})&-\mathsf{E}(\widehat{u}^{n+1})\\ &\geq-\int_{\Omega}\beta|\widehat{u}^{n+1}-u^{n}|^{4}\,\mathsf{d}\bm{\mathsf{x}}+\left(\nicefrac{{1}}{{\tau_{n}}}-\nicefrac{{1}}{{2}}\right)a_{0}(\widehat{u}^{n+1}-u^{n},\widehat{u}^{n+1}-u^{n}).\end{split}$ (55)
  3. Step 3:
    Based on (54) and (55), it can be shown by an inductive argument that there exists $0<\tau_{\max}\leq\nicefrac{{2}}{{5}}$ such that the energy decays. Indeed,
    $\mathsf{E}(u^{n})-\mathsf{E}(\widehat{u}^{n+1})\geq ca_{0}(\widehat{u}^{n+1}-u^{n},\widehat{u}^{n+1}-u^{n})$ (56)
    for some $c>0$ independent of $n$, but depending on $\tau_{\max}$. Thereby,
    $\mathsf{E}(u^{n+1})\leq\mathsf{E}(\widetilde{u}^{n+1})\leq\mathsf{E}(u^{n}),$ (57)
    where the first inequality follows from the mass growth obtained in the first step; this proves the first part of (a). Concerning (40), we simply refer to the proof in [29].
  4. Step 4:
    By the energy decay (57) and the positivity of $\mathsf{E}$, it immediately follows that $\mathsf{E}(u^{n})$ has some limit $\mathsf{E}^{\star}$. In turn, one can show that $u^{\star}$ is uniformly bounded in $\mathbb{H}$, and thus has a weakly convergent subsequence $\{u^{n_{j}}\}_{j}$ with limit $u^{\star}$; thanks to the Rellich-Kondrachov theorem, the convergence is strong in $\mathrm{L}^{p}(\Omega;\mathbb{C})$, for $61\leq p<$ 6. Moreover, recalling (56), the convergence of the energy further implies that $\left\|\widehat{u}^{n+1}-u^{n}\right\|_{\mathbb{H}}\to 0$ as $n\to\infty.$ Those convergence properties are key for all the following steps.
  5. Step 5:
    Based on step 4, it can further be shown that
    $(\mathsf{G}(u^{n_{j}}),u^{n_{j}})_{\mathrm{L}^{2}(\Omega)}\to(\mathsf{G}(u^{\star}),u^{\star})_{\mathrm{L}^{2}(\Omega)}=:(\lambda^{\star})^{-1}\quad\text{as}\ j\to\infty,$
    and, subsequently, we have $a_{u^{\star}}(u^{\star},v)=\lambda^{\star}(u^{\star},v)$ for all $v\in\mathbb{H}.$ In particular, the claim in (c) is verified.
  6. Step 6:
    A lengthy calculation discloses that $\mathsf{E}(u^{\star})=\mathsf{E}^{\star}$.
  7. Step 7:
    It remains to verify that the subsequence $\{u^{n_{j}}\}_{j}$ converges to $u^{\star}$ strongly in $\mathbb{H}$. Since the proof of this claim in the present work is somewhat more involved than in the case without the angular rotation momentum, we shall provide some details. First of all, recall that $\mathsf{E}(u^{n_{j}})\to\mathsf{E}(u^{\star})$, $u^{n_{j}}\to u^{\star}$ strongly in $L^{p}(\Omega)$, for $61\leq p<$ 6, and $u^{n_{j}}\hookrightarrow u^{\star}$ weakly in $\mathbb{H}$. Consequently, we find that
    $\lim_{j\to\infty}\left(\frac{1}{2}\left\|u^{n_{j}}\right\|_{\mathbb{H}}-\int_{\Omega}i\omega\overline{u}^{n_{j}}(\bm{\mathsf{A}}\cdot\nabla u^{n_{j}})\,\mathsf{d}\bm{\mathsf{x}}\right)=\frac{1}{2}\left\|u^{\star}\right\|_{\mathbb{H}}-\int_{\Omega}i\omega\overline{u^{\star}}(\bm{\mathsf{A}}\cdot\nabla u^{\star})\,\mathsf{d}\bm{\mathsf{x}}.$
    In particular, if we can show that
    00 0 $\leq\left|\int_{\Omega}\overline{u}^{n_{j}}(\bm{\mathsf{A}}\cdot\nabla u^{n_{j}})\,\mathsf{d}\bm{\mathsf{x}}-\int_{\Omega}\overline{u^{\star}}(\bm{\mathsf{A}}\cdot\nabla u^{\star})\,\mathsf{d}\bm{\mathsf{x}}\right|$
    $\leq\int_{\Omega}\left|u^{n_{j}}-u^{\star}\right|\left|\bm{\mathsf{A}}\cdot\nabla u^{n_{j}}\right|\,\mathsf{d}\bm{\mathsf{x}}+\left|\int_{\Omega}\left(\overline{u^{\star}}(\bm{\mathsf{A}}\cdot\nabla u^{n_{j}})\,\mathsf{d}\bm{\mathsf{x}}-\overline{u^{\star}}(\bm{\mathsf{A}}\cdot\nabla u^{\star})\right)\,\mathsf{d}\bm{\mathsf{x}}\right|$
    $=:I_{1}+I_{2}.$
    Concerning the first term, we recall that $\{u^{n_{j}}\}$ is a weakly convergent sequence in $\mathbb{H}$, and thus is uniformly bounded. Therefore, the Cauchy–Schwarz inequality yields that
    for some constant $C>0$ independent of $j$. To deal with the second term, we note that $v\mapsto\int_{\Omega}\overline{u^{\star}}(\bm{\mathsf{A}}\cdot\nabla v)\,\mathsf{d}\bm{\mathsf{x}}$ is a linear operator on $\mathbb{H}$, and thus $I_{2}\to 0$ as $j\to\infty$ thanks to the weak convergence of the sequence. Together with the first part, this concludes the proof of the strong convergence.

### v2

@@ 4.1.2. Non-linear GPE with optical lattice potential
Figure 3. Non-linear GPE with optical lattice potential. (a) Computed ground-state $u_{N}$; (b) Comparison of the error with respect to the square root of the number of degrees of freedom; (c) Final hphp-mesh.
In this second example, we consider a more challenging problem with an oscillating potential function $V$; to this end, we select $\Omega=(-6,6)^{2}$, $\beta=1000$, $V=\nicefrac{{1}}{{2}}(x^{2}+y^{2})+20+20\sin(2\pi x)\sin(2\pi y)$, and again $\omega=0$. The computed approximation to the ground-state is depicted in Figure 3(a); in this setting $u_{\mathrm{GS}}$ resembles a ‘hedgehog’ structure consisting of a large number of very thin spikes. Computations undertaken in [29] and [26] give values of the ground-state energy $\mathsf{E}(u_{\mathrm{GS}})\approx 30.40965$ and $\mathsf{E}(u_{\mathrm{GS}})\approx 30.387533$, respectively. Based on exploiting hphp–adaptivity, we have computed a reference value of $\mathsf{E}(u_{\mathrm{GS}})\approx\underline{30.3874145736}32611$.
The performance of the proposed hphp–refinement algorithm is depicted in Figure 3(b); here, again we observe exponential convergence of the error in the computed ground-state energy, despite the fine scale structures present in the ground-state. Moreover, hphp–refinement clearly outperforms its hh–version counterpart. The final hphp–mesh shown in Figure 3(c) indicates that some hh–refinement is necessary across most of the domain, with again pp–enrichment being undertaken in smooth regions.
@@ 4.1.3. Non-linear energy functional with a nonsymmetric potential $V$
Figure 4(b) again highlights the potential benefits of employing hphp–adaptivity: exponential convergence of the error in the computed ground-state energy is observed, as well as its superiority in terms of the computed error versus the number of degrees of freedom employed, when compared with mesh subdivision (hh–refinement). The final hphp–mesh depicted in Figure 4(c) highlights that some $h$-refinement is necessary around the top of the hill, while largely pp–enrichment is undertaken elsewhere.
@@
Figure 4. Non-linear GPE with a nonsymmetric potential $V$. (a) Computed ground-state $u_{N}$; (b) Comparison of the error with respect to the square root of the number of degrees of freedom; (c) Final hphp-mesh.
@@ 4.2. GPE with angular momentum operator $\omega>0$
Figure 5. GPE with angular momentum: $\omega=0.5$, $\beta=10$. (a) Computed ground-state density $|u_{N}|^{2}$; (b) Comparison of the error with respect to the square root of the number of degrees of freedom; (c) Final hphp-mesh.
Finally, in this section, we now turn our attention to the case when the angular momentum operator is present in the underlying energy functional, i.e., when $\omega>0$. To this end, we set $\Omega=(-L,L)$, for some $L>0$, $V=\nicefrac{{1}}{{2}}(x^{2}+y^{2})$, and consider the range of $\omega$ and $\beta$, for a given $L$, presented in Table 1. In addition, here we report the minimal energy computed using the proposed hphp–refinement algorithm based on selecting the initial condition
where
@@
Figure 6. GPE with angular momentum: $\omega=0.75$, $\beta=100$. (a) Computed ground-state density $|u_{N}|^{2}$; (b) Comparison of the error with respect to the square root of the number of degrees of freedom; (c) Final hphp-mesh.
Numerical experiments for each of the values of $\omega$ and $\beta$ given in Table 1 are presented in Figures 58. As before, we show the computed ground-state, the comparison between the error in the approximated ground-state energy when employing both hphp– and hh–refinement, and the final hphp–mesh. We note that as $\omega$ is increased, the number of vortices present in the ground-state increases; for the case when $\omega=0.9$ and $\beta=1000$, the ground-state features over 50 vortices arranged in a regular triangular lattice referred to as the Abrikosov lattice. In all cases we again observe that the proposed hphp–refinement algorithm leads to exponential convergence of the error in the approximated ground-state energy; moreover, the efficiency of employing hphp–refinement when compared to mesh subdivision alone is clearly highlighted. As the number of vortices increases with each of the test cases considered here, we observe that the accuracy attained for the computed ground-state energy decreases. Moreover, we also observe that more refinement of the computational mesh is necessary, before subsequent pp–refinement is employed to yield exponential convergence of the error in the ground-state energy, as $\omega$ is increased. In the simplest setting, i.e., when $\omega=0.5$ and $\beta=10$ we observe that the proposed hphp–refinement algorithm simply defaults to polynomial enrichment only.
@@
Figure 7. GPE with angular momentum: $\omega=0.8$, $\beta=500$. (a) Computed ground-state density $|u_{N}|^{2}$; (b) Comparison of the error with respect to the square root of the number of degrees of freedom; (c) Final hphp-mesh.
@@
Figure 8. GPE with angular momentum: $\omega=0.9$, $\beta=1000$. (a) Computed ground-state density $|u_{N}|^{2}$; (b) Comparison of the error with respect to the square root of the number of degrees of freedom; (c) Final hphp-mesh.
@@ 5. Conclusions
In this paper, we have considered the stationary non-linear Gross–Pitaevskii eigenvalue problem in the presence of a rotating magnetic field, which yields an intricate topology manifested by the presence of quantized vortices in the ground-state; the number and location of these vortices is unknown a priori. In particular, we have extended the energy-based adaptive finite element strategy presented in  [26] in two key directions: firstly, we have considered the case when a rotating magnetic field is present in the underlying energy functional, and secondly, we have generalised the adaptivity framework to allow for $hp$-refinements of the underlying discrete finite element approximation space. Towards these goals, we have also presented a result on the convergence of a discrete iteration scheme (Theorem 2.7) which is based on a continuous projected gradient flow that respects the energy-based topology of the underlying problem at hand. The key feature of the proposed energy-adaptive $hp$-mesh refinement strategy is that it is based solely on potential energy-decay which can be computed locally and does not require sophisticated localisation schemes as is standard for residual-based a posteriori error estimators. The resulting method thereby intertwines the $hp$-mesh refinement iterative procedure with the iterative non-linear solver for the underlying problem within a unified framework in the sense that the numerical approximation to the non-linear problem is computed in tandem while the $hp$-mesh is adaptively refined. The performance of this proposed strategy is demonstrated on a diverse set of benchmark test problems that highlight the exponential convergence of the error in the computed ground-state energy with respect to the number of degrees of freedom in the resulting $hp$-finite element space.
@@ Acknowledgments
T. P. W. acknowledges the financial support of the Swiss National Science Foundation (SNF), Grant No. $200021\underline{\phantom{-}}212868$. B. S. acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), Project No. 442047500, through the Collaborative Research Center “Sparsity and Singular Structures” (SFB 1481). Finally, we thank Eric Cancès for encouraging discussions and motivating us to extend the framework to the case with rotating magnetic field.